# Recursive Gaussian Process State Space Model

### v3

@@ 7.1 Score for Selection of Inducing Point
Firstly, for $D_{2}$, the log-term in the intergration is:
@@
where $\bm{Q}_{dd}$ denotes the $i_{\mathrm{d}}$th diagonal element of matrix $\bm{Q}$, and $\bm{Q}_{dl}$ denotes the $i_{\mathrm{d}}$th row of $\bm{Q}$ with $\bm{Q}_{dd}$ removed.
Using (21), (38) and (39), the $u_{\mathrm{d}}$ relevant term in $D_{2}$ is:
@@
where $p(\bm{a})$ is a $n_{a}$-dimension Gaussian distribution with covariance $\bm{\Sigma}_{a}$. Deonte the covariance of joint distribution $q(\bm{X}_{t})$ that marginal out $u_{\mathrm{d}}$ as $\bm{\Sigma}_{t,l}$, and deonte the inversion of orignial covariance as $\bm{\Omega}=\bm{\Sigma}_{t}^{-1}$. Then, given (44), the $u_{\mathrm{d}}$ relevant term in $D_{3}$ is:
@@
where $\bm{\Omega}_{dd}$ is the element of the matrix $\bm{\Omega}$ corresponding to the discarded point $u_{\mathrm{d}}$, namely, its $(d_{x}+i_{\mathrm{d}})$-th diagonal element. In addition, (45) is dervied by using the properties of the determinants of block matrices and the matrix inversion formula (see Appendix A of [52]).
Based on the KL divergence (21) and combing (40), (43) and (45), we have the scroe:
@@
7.2 Prediction Equation without Adding Points
@@
7.3 Derivation Details of GP Hyperparameters Optimization
This section derives the moments of the approximate joint distribution $q_{new}(\bm{x}_{t},\bm{f})$ in (29), along with the optimizing objective for GP hyperparameters.
Firstly, to derive the moments of the approximate joint distribution $q_{new}(\bm{x}_{t},\bm{f})$, we can utilize the following result:
$\dfrac{p(\bm{u};\bm{\theta}_{new})}{p(\bm{u};\bm{\theta}_{old})}$ $=\dfrac{\mathcal{N}(\bm{u}|\bm{0},\bm{K}_{uu}^{new})}{\mathcal{N}(\bm{u}|\bm{0},\bm{K}_{uu}^{old})}$ (51)
@@ 7.3 Derivation Details of GP Hyperparameters Optimization
$\Delta\bm{K}$ $=[(\bm{K}_{uu}^{new})^{-1}-(\bm{K}_{uu}^{old})^{-1}]^{-1}$ (52)
$Z$ $\propto\dfrac{|\bm{K}_{uu}^{old}|^{1/2}|\Delta\bm{K}|^{1/2}}{|\bm{K}_{uu}^{new}|^{1/2}}$
@@
$q_{new}(\bm{x}_{t},\bm{u})\propto\mathcal{N}(\bm{0}|\bm{u},\Delta\bm{K})q_{old}(\bm{x}_{t},\bm{u})$ (53)
whose moments can be obtained by the correction eqution of Kalman filter:
@@
$\tilde{\bm{G}}=\bm{\Sigma}_{t}^{old}\tilde{\bm{H}}^{T}\left(\bm{S}_{uu}^{old}+\Delta\bm{K}\right)^{-1}$
$\bm{\xi}_{t}^{new}=\bm{\xi}_{t}^{old}-\tilde{\bm{G}}\bm{\xi}_{t}^{old}$
$\bm{\Sigma}_{t}^{new}=\bm{\Sigma}_{t}^{old}-\tilde{\bm{G}}\tilde{\bm{H}}\bm{\Sigma}_{t}^{old}$
@@
$\log\dfrac{p(\bm{y}_{1:t};\bm{\theta}_{new})}{p(\bm{y}_{1:t};\bm{\theta}_{old})}$ (55)
$=\log\int Z\mathcal{N}(\bm{0}|\bm{u},\Delta\bm{K})q_{old}(\bm{x}_{t},\bm{u})\,\mathrm{d}\bm{x}_{t}\,\mathrm{d}\bm{u}$
$=\log\mathcal{N}(\bm{0}|\bm{m}_{u}^{old},\Delta\bm{K}+\bm{S}_{uu}^{old})+\log Z$
Then, by dropping some $\bm{\theta}_{new}$-irrelevant terms, the optimization objective can be transformed to minimize:
@@
$\mathcal{L}_{1}$ $=(\bm{m}_{u}^{old})^{T}(\bm{S}_{uu}^{old}+\Delta\bm{K})^{-1}\bm{m}_{u}^{old}$ (57)
$\mathcal{L}_{2}$ $=-\log|\Delta\bm{K}|+\log|\bm{K}_{uu}^{new}|+\log|\Delta\bm{K}+\bm{S}_{uu}^{old}|$
$=\log|\bm{K}_{uu}^{new}|+\log|\bm{I}_{n_{u}}+\Delta\bm{K}^{-1}\bm{S}_{uu}^{old}|$
$=\log|\bm{K}_{uu}^{new}+[\bm{I}_{n_{u}}-\bm{K}_{uu}^{new}(\bm{K}_{uu}^{old})^{-1}]\bm{S}_{uu}^{old}|$
Here, we avoid the appearance of $\Delta\bm{K}$ in the loss function through some operations, because when $\bm{K}_{uu}^{new}=\bm{K}_{uu}^{old}$, $\Delta\bm{K}$ is not well-defined.
7.4 Stable Implementation Method
@@
$\bm{\Sigma}_{t+1}^{-}$ $=\begin{bmatrix}\bm{\Phi}_{t}\bm{L}_{t}&\bm{D}_{f}\\ \end{bmatrix}\begin{bmatrix}\bm{L}_{t}^{T}\bm{\Phi}_{t}^{T}\\ \bm{D}_{f}^{T}\\ \end{bmatrix}$ (58)
where $\bm{D}_{f}$ is the Cholesky factor of the process noise covariance $\bm{\Sigma}_{p,\bm{X}}$. Then, through the following QR decomposition:
$\begin{bmatrix}\bm{L}_{t}^{T}\bm{\Phi}_{t}^{T}\\ \bm{D}_{f}^{T}\\ \end{bmatrix}=\bm{O}\bm{R}$ (59)
@@
For hyperparameter optimization, since $\Delta\bm{K}$ in (32) may not be positive definite as illustrated in Appendix 7.3, we cannot update the Cholesky factor as in the correction step. Therefore, the updated Cholesky factor is evaluated by directly decomposing the updated covariance $\bm{\Sigma}_{t}^{\text{new}}$, which is obtained using (32):

### v4

@@ Appendix A Score for Selection of Inducing Point
Firstly, for $D_{2}$, the log-term in the integral is:
@@
where $\bm{Q}_{dd}$ denotes the $i_{\mathrm{d}}$-th diagonal element of matrix $\bm{Q}$, and $\bm{Q}_{dl}$ denotes the $i_{\mathrm{d}}$-th row of $\bm{Q}$ with $\bm{Q}_{dd}$ removed.
Using (21), (38) and (39), the $u_{\mathrm{d}}$-relevant term in $D_{2}$ is:
@@
where $p(\bm{a})$ is a $n_{a}$-dimension Gaussian distribution with covariance $\bm{\Sigma}_{a}$. Deonte the covariance of joint distribution $q(\bm{X}_{t})$ that marginal out $u_{\mathrm{d}}$ as $\bm{\Sigma}_{t,l}$, and denote the inversion of original covariance as $\bm{\Omega}=\bm{\Sigma}_{t}^{-1}$. Then, given (44), the $u_{\mathrm{d}}$ relevant term in $D_{3}$ is:
@@
where $\bm{\Omega}_{dd}$ is the element of the matrix $\bm{\Omega}$ corresponding to the discarded point $u_{\mathrm{d}}$, namely, its $(d_{x}+i_{\mathrm{d}})$-th diagonal element. In addition, (45) is derived by using the properties of the determinants of block matrices and the matrix inversion formula (see Appendix A of [39]).
Based on the KL divergence (21) and combing (40), (43) and (45), we have the score:
@@
Appendix B Prediction Equation without Adding Points
@@
Appendix C Derivation Details of GP Hyperparameters Optimization
This section derives the moments of the approximate joint distribution $q_{\mathrm{new}}(\bm{x}_{t},\bm{f})$ in (29), along with the optimizing objective for GP hyperparameters.
Firstly, to derive the moments of the approximate joint distribution $q_{\mathrm{new}}(\bm{x}_{t},\bm{f})$, we can utilize the following result:
$\dfrac{p(\bm{u};\bm{\theta}_{\mathrm{new}})}{p(\bm{u};\bm{\theta}_{\mathrm{old}})}$ $=\dfrac{\mathcal{N}(\bm{u}|\bm{0},\bm{K}_{uu}^{\mathrm{new}})}{\mathcal{N}(\bm{u}|\bm{0},\bm{K}_{uu}^{\mathrm{old}})}$ (51)
@@ Appendix C Derivation Details of GP Hyperparameters Optimization
$\Delta\bm{K}$ $=[(\bm{K}_{uu}^{\mathrm{new}})^{-1}-(\bm{K}_{uu}^{\mathrm{old}})^{-1}]^{-1}$ (52)
$Z$ $\propto\dfrac{|\bm{K}_{uu}^{\mathrm{old}}|^{1/2}|\Delta\bm{K}|^{1/2}}{|\bm{K}_{uu}^{\mathrm{new}}|^{1/2}}$
@@
$q_{\mathrm{new}}(\bm{x}_{t},\bm{u})\propto\mathcal{N}(\bm{0}|\bm{u},\Delta\bm{K})q_{\mathrm{old}}(\bm{x}_{t},\bm{u})$ (53)
whose moments can be obtained by the correction equation of Kalman filter:
@@
$\log\dfrac{p(\bm{y}_{1:t};\bm{\theta}_{\mathrm{new}})}{p(\bm{y}_{1:t};\bm{\theta}_{\mathrm{old}})}$ (55)
$=\log\int Z\mathcal{N}(\bm{0}|\bm{u},\Delta\bm{K})q_{\mathrm{old}}(\bm{x}_{t},\bm{u})\,\mathrm{d}\bm{x}_{t}\,\mathrm{d}\bm{u}$
$=\log\mathcal{N}(\bm{0}|\bm{m}_{u}^{\text{old}},\Delta\bm{K}+\bm{S}_{uu}^{\text{old}})+\log Z$
Then, by dropping some $\bm{\theta}_{\mathrm{new}}$-irrelevant terms, the optimization objective can be transformed to minimize:
@@
$\mathcal{L}_{1}$ $=(\bm{m}_{u}^{\text{old}})^{T}(\bm{S}_{uu}^{\text{old}}+\Delta\bm{K})^{-1}\bm{m}_{u}^{\text{old}}$ (57)
$\mathcal{L}_{2}$ $=-\log|\Delta\bm{K}|+\log|\bm{K}_{uu}^{\mathrm{new}}|+\log|\Delta\bm{K}+\bm{S}_{uu}^{\text{old}}|$
$=\log|\bm{K}_{uu}^{\mathrm{new}}|+\log|\bm{I}_{n_{u}}+\Delta\bm{K}^{-1}\bm{S}_{uu}^{\text{old}}|$
$=\log|\bm{K}_{uu}^{\mathrm{new}}+[\bm{I}_{n_{u}}-\bm{K}_{uu}^{\mathrm{new}}(\bm{K}_{uu}^{\mathrm{old}})^{-1}]\bm{S}_{uu}^{\text{old}}|$
Here, we avoid the appearance of $\Delta\bm{K}$ in the loss function through some operations, because when $\bm{K}_{uu}^{\mathrm{new}}=\bm{K}_{uu}^{\mathrm{old}}$, $\Delta\bm{K}$ is not well-defined.
Appendix D Stable Implementation Method
@@
$\bm{\Sigma}_{t+1}^{-}$ $=\begin{bmatrix}\bm{\Phi}_{t}\bm{L}_{t}&\bm{D}_{p}\\ \end{bmatrix}\begin{bmatrix}\bm{L}_{t}^{T}\bm{\Phi}_{t}^{T}\\ \bm{D}_{p}^{T}\\ \end{bmatrix}$ (58)
where $\bm{D}_{p}$ is the Cholesky factor of the process noise covariance $\bm{\Sigma}_{p,\bm{X}}$. Then, through the following QR decomposition:
$\begin{bmatrix}\bm{L}_{t}^{T}\bm{\Phi}_{t}^{T}\\ \bm{D}_{p}^{T}\\ \end{bmatrix}=\bm{O}\bm{R}$ (59)
@@
For hyperparameter optimization, since $\Delta\bm{K}$ in (30) may not be positive definite as illustrated in Appendix C, we cannot update the Cholesky factor as in the correction step. Therefore, the updated Cholesky factor is evaluated by directly decomposing the updated covariance $\bm{\Sigma}_{t}^{\text{new}}$, which is obtained using (30):